Find the equation of the line that passes through the equation point P and is Parallel to the line L.
P(2,1), L:3x+y=-2
I'm so lost.

Answer by Edwin McCravy(20055)

Find the equation of the line that passes through the equation point P and is Parallel to the line L.
P(2,1), L:3x+y=-2

We will graph the given line by first solving for y:

3x + y = -2
     y = -3x - 2

 x |  y
 2 | -8
-2 |  4

So we draw the graph by plotting those two points:



and drawing a green line through them:



Next we plot the point P(2,1)



We compare the equation

     y = -3x - 2
to
      y = mx + b

and observe that m = -3

Lines that are parallel have the same slope.

We now use the point-slope form of a line

y - y1 = m(x - x1)

and substitute m = -3 and (x1,y1) = (2,1)

y - 1 = -3(x - 2)

y - 1 = -3x + 6
    
    y = -3x + 7

That's the answer but let's check by graphing it.
Get a couple points on it:
 
   y = -3x + 7

 x |  y
 1 |  4
 4 | -5




Draw a red line through them:



Yes, the red line looks parallel to the green line, and it
appears to go through P(2,1).  So the answer is 

   y = -3x + 7

Edwin